Justin has the utility function U = xy, with the marginal utilities MUx = y and MUy = x. The price of x is $2, the price of y is py, and his income is 40. When he maximizes utility subject to his budget constraint, he purchases 5 units of y.

(a)   What must be the price of y and the amount of x consumed?

(b)  Prove that this allocation follows the equi-marginal principle

What would be the new bundles of x, y if Px was $3


1
Expert's answer
2021-05-16T19:20:57-0400

a) "Utility\ function = U=xy"

"MU_x=y"

"MU_y=x"

"P_x= \$2"

"P_y=py"

Income= $40.

Budget line will be "M=P_xx+P_yy"

"40=2x+5y"

"\frac{MU_x}{P_x}=\frac{MU_y}{P_y}"

"\frac{y}{2}=\frac{x}{5}"

"5y=2x"

"y=\frac{2}{5}x"

"40=2x+5(\frac{2}{5}x)"

"40=4x"

"x=10"

Amount of x consumed will be 10 units.

"y=\frac{2}{5}\times 10"

"y=4"

The price of Y will therefore, be $4.


b) In equi-marginal principle, "MU_x=P_x"

Therefore, "y=P_x"

"y=4" and "P_x=4". This proves that the allocation above follows the equi-marginal principle.


c) if "P_x=3"

"\frac{y}{3}=\frac{x}{5}"

"5y=3x"

"y=\frac{3}{5}x"

"40=2x+5(\frac{3}{5}x)"

"40=5x"

"x=8"

"y=\frac{3}{5}\times 8"

"y=\frac{24}{5}"